Machinery was purchased on January 1 for $96,040.00. The machinery has an estimated life of seven years and an estimated salvage value of $9,000. Double-declining-balance depreciation for the second year would be (round calculations to the nearest dollar):

a.$18,600

b.$19,100

c.$19,600

d.$20,600
Answer #1

Answer

  • Correct Answer = Option ‘C’ $ 19,600
  • Working

A

Cost

$            96,040.00

B

Residual Value

$              9,000.00

C=A - B

Depreciable base

$            87,040.00

D

Life [in years]

7

E=C/D

Annual SLM depreciation

$            12,434.29

F=E/C

SLM Rate

14.29%

G=F x 2

DDB Rate

28.57%

Year

Beginning Book Value

Depreciation rate

Depreciation expense

Ending Book Value

1

$              96,040.00

28.57%

$        27,440.00

$             68,600.00

2

$              68,600.00

28.57%

$         19,600.00 Answer

$             49,000.00